OpenStudy (anonymous):

Quick question! How would I FOIL (3x + 7y)^2? I'm a little confused with the x's and y's...

OpenStudy (phi):

you write it as (3x+7y)(3x+7y) F first: 3x*3x = ?

OpenStudy (anonymous):

3x^2?

OpenStudy (phi):

remember, math people don't like to type 3x*3x is the same as 3*x*3*x and when you multiply, you can change the order, so it could also be written 3*3*x*x can you simplify it now?

OpenStudy (anonymous):

so 9x^2?

OpenStudy (phi):

yes now O outer: 3x*7y (or for you, 3*x*7*y )

OpenStudy (phi):

yes, it is very possible x*y is xy (because xy means x*y) now I inner: 7y*3x (same thing as before)

OpenStudy (anonymous):

21xy

OpenStudy (phi):

and L last: 7y*7y

OpenStudy (anonymous):

49y^2

OpenStudy (anonymous):

Would I combine them to be 42x^2y^2?

OpenStudy (phi):

so we have 9x^2 + 21xy + 21xy + 49y^2 the xy terms can be combined: if you factor out xy from (21xy + 21xy)= (21+21)*xy= 42*xy or just 42xy 9x^2 +42xy + 49y^2 or you can say I have 21 xy's (whatever that is) and add 21 more xy's to get 42 xy's

OpenStudy (phi):

Here is the idea for combining "like terms" a "like term" has exactly the same letters (to the same power) so x^2 matches x^2 but x^2 does NOT match x or x^3 (different powers) and x does NOT match y but xy matches yx (order does not matter) so xy^2 matches xy^2 but not x^2y

OpenStudy (phi):

If you want more info, maybe this helps http://www.khanacademy.org/math/algebra/polynomials/v/multiplying-binomials

OpenStudy (anonymous):

okay, so the answer would just be 3x^2 + 42xy + 49y^2?

OpenStudy (anonymous):

I mean 9x^2****

OpenStudy (anonymous):